Show by dimensional analysis that the product of resistance and capacitance has the dimension of time.


1
Expert's answer
2022-04-29T12:30:44-0400
"[\Omega]=\rm[\frac{kg\cdot m^2}{s^3\cdot A^2}]""[F]=\rm[\frac{s^4\cdot A^2}{kg\cdot m^2}]""[RC]=[\Omega F]=\rm[\frac{kg\cdot m^2}{s^3\cdot A^2}*\frac{s^4\cdot A^2}{kg\cdot m^2}]=[s]"